OpenStudy (anonymous):

Simplify the expression 2-(4-t) when t = 1

OpenStudy (anonymous):

2-4+t=-2+t

OpenStudy (anonymous):

so if t = 1 -2+1 = -1 therefore -1

OpenStudy (ashleyisakitty):

2-(4-t) Replace t for 1. 2-(4-1). Now you have to remember the order of operations. The first step is to figure out whats in the parenthesis. 4-1. 4-1=3. Now we have 2-(3). That is your simplified equation. As Mathmetics said, the answer is -1.